OpenStudy (anonymous):

An airplane is heading due south at a speed of 600 km h. If a wind begins blowing from the southwest at a speed of 100 km h (average), calculate: (a) the velocity (magnitude and direction) of the plane relative to the groun

OpenStudy (anonymous):

Please keep in mind the wind is blowing FROM the southwesterly direction, not towards.

OpenStudy (anonymous):

You're 100% correct. Let me fix the answer :D

OpenStudy (anonymous):

Use pythagorean theorem for speed: \[x^2+y^2=s^2\] (Since the wind is moving from the southwest, you need to subtract 100/sqrt(2) km/hr to the south direction and 100/sqrt(2)km/hr to the west direction) \[s^2=(100/(√2))^2+(600-100/(√2))^2\] \[s=533.991747\] For the direction you need to get the angle: \[tan(α)=(−600+100/(√2)/−100/(√2))\] \[α=262.390598\]

OpenStudy (anonymous):

Check the math, but the equations are correct :D

OpenStudy (anonymous):

Thanks a bunch. That's exactly what I got. Boy, do I hate book errors. XD

OpenStudy (anonymous):

Me too...